OpenStudy (anonymous):

A television set shoots out a beam of electrons. The beam current is 10uA.How many electrons strike the TV screen each second? How much charge strike the screen in a minute

OpenStudy (anonymous):

10uA is 10 microamps or \[10^{-5}\] amps. 1 Amp is \[6.241 \times 10^{18}\] electrons passing by a section per second. So 10uA will be:\[\frac{6.241 \times 10^{18}}{10^{5}} = 6.241 \times 10^{13}\] electrons per second. Which is interesting because it doesn't depend on the cross sectional that the electrons travel through due to the way current is defined. Which I find strange. As for charge well 1 amp is a coulomb every second. A coulomb, C, is the unit of charge. Since we have one hundred thousandth of an amp and one amp is one coulomb per second, it makes sense that we have one hundred thousandth of a colomb every second. Times this by your amount of seconds, 60, gives \[10 ^{-5} C \times 60 = 6 \times 10 ^{-4}C\] which is 600 uC , or 600 millicoulomb's of charge in a minute